OpenStudy (anonymous):

A state's license plates consist of four letters followed by three numerals, and 245 letter arrangements are not allowed. How many plates can the state issue?

OpenStudy (anonymous):

four letters \(26^4\) the numerals \(10^3=1000\) so by the counting principle \[26^4\times 10^3\] but i guess you have to subtract something

OpenStudy (anonymous):

need to subtract \(245,000\) to account for the \(245\) not permissible letter arrangements

OpenStudy (anonymous):

but subtract what number is the answer 1000

OpenStudy (anonymous):

no not 1000

OpenStudy (anonymous):

You have: \[26^{3} \times 10^{3} = 17,576,000\]different ways but 249 ways are not allowed. So you really have \[17,576,000-249 = 17,575,751 \]different ways.

OpenStudy (anonymous):

number of possible letter arrangements is actually \[26^4-245\]

OpenStudy (anonymous):

@Rachella i think you need to subtract more

OpenStudy (anonymous):

Okay but with the same numbers because am not sure

OpenStudy (anonymous):

and 245 letter arrangements are not allowed

OpenStudy (anonymous):

number of possible letter arrangements is \[26^4-245\] number of total arrangements therefore is \[(26^4-245)\times 1000\]

OpenStudy (anonymous):

Oh let me see what I get

OpenStudy (anonymous):

For some reason it not allowing me to get the answer on the calculator

OpenStudy (anonymous):

I got an answer of 456731000 is that right